# A Comprehensive Survey on the Trustworthiness of Large Language Models in Healthcare

### v1

@@ Abstract
The application of large language models (LLMs) in healthcare has the potential to revolutionize clinical decision-making, medical research, and patient care. As LLMs are increasingly integrated into healthcare systems, several critical challenges must be addressed to ensure their reliable and ethical deployment. These challenges include truthfulness, where models generate misleading information; privacy, with risks of unintentional data retention; robustness, requiring defenses against adversarial attacks; fairness, addressing biases in clinical outcomes; explainability, ensuring transparent decision-making; and safety, mitigating risks of misinformation and medical errors. Recently, researchers have begun developing benchmarks and evaluation frameworks to systematically assess the trustworthiness of LLMs. However, the trustworthiness of LLMs in healthcare remains underexplored, lacking a systematic review that provides a comprehensive understanding and future insights into this area. This survey bridges this gap by providing a comprehensive overview of the recent research of existing methodologies and solutions aimed at mitigating the above risks in healthcare. By focusing on key trustworthiness dimensions including truthfulness, privacy and safety, robustness, fairness and bias, and explainability, we present a thorough analysis of how these issues impact the reliability and ethical use of LLMs in healthcare. This paper highlights ongoing efforts and offers insights into future research directions to ensure the safe and trustworthy deployment of LLMs in healthcare.
@@ A Comprehensive Survey on the Trustworthiness of Large Language Models in Healthcare
Manar Aljohani1, Jun Hou1, Sindhura Kommu1, Xuan Wang1, 1Department of Computer Science, Virginia Tech, Blacksburg, VA, USA (manara, junh, sindhura, xuanw)@vt.edu
@@ 1 Introduction
Truthfulness, defined as "the accurate representation of information, facts, and results by an AI system" Huang et al. (2024), is critical in healthcare, as inaccuracies can lead to misdiagnoses or inappropriate treatment recommendations. Ensuring that generated information is both accurate and aligned with verified medical knowledge is essential. Additionally, privacy concerns arise from the risk of exposing sensitive patient data during model training and usage, potentially leading to breaches or violations of regulations such as HIPAA (Health Insurance Portability and Accountability Act) and GDPR (General Data Protection Regulation). Ensuring patient confidentiality while leveraging LLMs for diagnostics and treatment recommendations is a critical challenge. Safety, defined as “ensuring that LLMs do not answer questions that can harm patients or healthcare providers in healthcare settings” Han et al. (2024b), further underscores the necessity of implementing stringent safeguards to mitigate harm. Robustness refers to an LLM’s ability to consistently generate accurate, reliable, and unbiased outputs across diverse clinical scenarios while minimizing errors, hallucinations, and biases. It also encompasses the model’s resilience against adversarial attacks, ensuring that external manipulations do not compromise its integrity. A truly robust LLM in healthcare must demonstrate stability, reliability, and fairness, even when faced with noisy, ambiguous, or adversarial inputs, thereby safeguarding patient safety and supporting clinical decision-making. Similarly, fairness and bias must be addressed to prevent discriminatory patterns in model predictions, which could lead to unequal treatment recommendations and exacerbate healthcare disparities. Furthermore, the explainability of LLMs, which ensures that model outputs are interpretable and transparent, plays a vital role in fostering trust and allowing informed decision-making by healthcare professionals. The lack of transparency in model reasoning complicates clinical adoption and raises concerns about accountability.
@@
Tackling these challenges is essential for the trustworthy and ethical implementation of LLMs in healthcare. Recently, researchers have begun developing benchmarks and evaluation frameworks to systematically assess the trustworthiness of LLMs Huang et al. (2024). The trustworthiness of LLMs in healthcare is gaining increasing attention due to its significant social impact. However, there is currently no systematic review that provides a comprehensive understanding and future insights into this area. To bridge this gap, we present a comprehensive survey that explores these trust-related dimensions in detail, reviewing existing datasets, solutions, and methodologies aimed at improving the trustworthiness of LLMs in healthcare.
@@ 2 Datasets, Models, and Tasks
We first conducted an extensive search for papers on the trustworthiness of LLMs in healthcare. Our search utilized a range of keyword combinations, including terms such as ‘large language models,’ ‘foundation model,’ ‘medical,’ ‘clinical,’ ‘explainability,’ ‘truthfulness,’ ‘trustworthiness,’ ‘safety,’ ‘fairness,’ ‘robustness,’ and ‘privacy.’ We explored several reputable venues, including Arxiv, PubMed, ACL, EMNLP, NAACL, ICML, NeurIPS, ICLR, KDD, Nature, Science, AAAI, and IJCAI, with a focus on recent publications post-2021. After reviewing the search results, we identified a total of 30,595 papers. Following the removal of duplicates, we narrowed the focus to 69 papers that specifically addressed the truthfulness, privacy, safety, robustness, fairness, bias, and explainability of LLMs in the healthcare domain.
@@
We then summarized all the datasets, models, and tasks relevant to research on trust in LLMs for healthcare, providing a comprehensive overview of their applications and contributions to this domain. The datasets used in studies of trust in LLMs for healthcare are categorized by the dimensions of trustworthiness they address in Appendix A, where we highlight key details such as data type, content, task, and dimensions of trustworthiness. The content of each dataset specifies its composition, while the task refers to the primary purpose for which the dataset is utilized. The data type varies across datasets and includes web-scraped data, curated domain-specific datasets, public text corpora, synthetic data, real-world data, and private datasets, providing a comprehensive overview of their relevance to healthcare applications. The models assessed in studies on trust in LLMs for the healthcare domain are outlined, along with their trustworthiness dimensions, in Appendix B, where we summarized key details such as the model name, release year, task, and the institution responsible for its development. The tasks covered various primary focuses of LLMs in healthcare. Based on insights from the survey by Liu et al. (2024), these tasks are outlined as follows:
@@
{forest}
forked edges, for tree= grow=east, reversed=true, anchor=base west, parent anchor=east, child anchor=west, base=center, font=, rectangle, draw=hidden-draw, rounded corners, align=left, text centered, minimum width=4em, edge+=darkgray, line width=1pt, s sep=3pt, inner xsep=2pt, inner ysep=3pt, line width=0.8pt, ver/.style=rotate=90, child anchor=north, parent anchor=south, anchor=center, , where level=1text width=14em,font=,, where level=2text width=18em,font=,, where level=3text width=18em,font=,, [ Trustworthiness of LLMs in Healthcare, ver [ Truthfulness, fill=blue!10 [ Med-HALT Pal et al. (2023), Med-HVL Yan et al. (2024b), Med-HallMark Chen et al. (2024), Semantic Entropy Farquhar et al. (2024),
SEPs Han et al. (2024a), Survey Ahmad et al. (2023), Self Reflection Ji et al. (2023), Model-Agnostic Hallucination Post-Processing Li et al. (2024),
Faithful Reasoning Tan et al. (2024), PubHealthTab Akhtar et al. (2022), HEALTHVER Sarrouti et al. (2021), CRITIC Gou et al. (2024),
Cross-Examination Cohen et al. (2023), KnowledgeEditor De Cao et al. (2021), leaf, text width=66em ] ] [ Privacy, fill=blue!10 [ Survey Das et al. (2024), Survey Yan et al. (2024a), Privacy Risks Pan et al. (2020), Federated Learning Zhao et al. (2024),
Differential Privacy Singh et al. (2024), SecureSQL Song et al. (2024), Memorize Fine-tuning Data Yang et al. (2024a),
clinical Note De-identification Altalla’ et al. (2025), De-identification Liu et al. (2023b), leaf, text width=66em ] ] [ Safety, fill=blue!10 [ Med-harm Han et al. (2024c), Medsafetybench Han et al. (2024b), UNIWIZ Das and Srihari (2024), Misinformation Attacks Han et al. (2024d),
MEDIC Kanithi et al. (2024), leaf, text width=66em ] ] [ Robustness, fill=blue!10 [ Survey Yuan et al. (2023), Survey Alberts et al. (2023), Stumbling Blocks Wang et al. (2024), LLM-TTA O’Brien et al. (2024), MedFuzz Ness et al. (2024),
Detecting Anomalies Rahman et al. (2024), Instruction Phrasings Ceballos-Arroyo et al. (2024), Adversarial Attacks Yang et al. (2024b),
Secure Your Model Tang et al. (2024), leaf, text width=66em ] ] [ Fairness and Bias, fill=blue!10 [ Evaluation StudyZack et al. (2024), BiasMedQA Schmidgall et al. (2024), Survey Parray et al. (2023), Instruction Fine-tuning Singhal et al. (2023),
Hurtful Words Zhang et al. (2020), Race-based Medicine Omiye et al. (2023), Detect Debunked Stereotypes Swaminathan et al. (2024),
EquityMedQA Pfohl et al. (2024), Superficial Fairness Alignment Wei et al. (2024), Examines Biased AI Adam et al. (2022),
Identify Biases Yang et al. (2024c), Quantifying Cognitive Biases Lin and Ng (2023), Mitigate Cognitive Biases Ke et al. (2024), CI4MRC Zhu et al. (2023), leaf, text width=66em ] ] [ Explanability, fill=blue!10 [ Knowledge Graphs Shariatmadari et al. (2024), LLMs and Explainable ML Elsborg and Salvatore (2023), Medical Imaging Explainability Ghosh et al. (2023),
MedViLaM Xu et al. (2024), MedThink Gai et al. (2024), MedExQA Kim et al. (2024), Causal Graphs Meet Thoughts Luo et al. (2025),
Retrieval and Reasoning on KGs Ji et al. (2024), TOSRR Liu et al. (2025), DDCoT Zheng et al. (2023), Layered Chain-of-Thought Prompting Sanwal (2025),
A ChatGPT Aided Explainable Framework Liu et al. (2023a), leaf, text width=66em ] ] ]
@@ 3 Trustworthiness of LLMs in Healthcare
We examine the challenges related to the trustworthiness of LLMs in healthcare, outlining key strategies for identifying and mitigating these concerns. From our literature review screening, we identified truthfulness, privacy and safety, robustness, fairness and bias, and explainability as key trustworthiness dimensions of LLMs as highlighted in TrustLLM Huang et al. (2024), particularly in healthcare. Figure 1 provides a summary of the recent research on trust in LLMs for healthcare across key dimensions of trustworthiness.

### v2

@@ Abstract
The application of large language models (LLMs) in healthcare holds significant promise for enhancing clinical decision-making, medical research, and patient care. However, their integration into real-world clinical settings raises critical concerns around trustworthiness, particularly around dimensions of truthfulness, privacy, safety, robustness, fairness, and explainability. These dimensions are essential for ensuring that LLMs generate reliable, unbiased, and ethically sound outputs. While researchers have recently begun developing benchmarks and evaluation frameworks to assess LLM trustworthiness, the trustworthiness of LLMs in healthcare remains underexplored, lacking a systematic review that provides a comprehensive understanding and future insights. This survey addresses that gap by providing a comprehensive review of current methodologies and solutions aimed at mitigating risks across key trust dimensions. We analyze how each dimension affects the reliability and ethical deployment of healthcare LLMs, synthesize ongoing research efforts, and identify critical gaps in existing approaches. We also identify emerging challenges posed by evolving paradigms, such as multi-agent collaboration, multi-modal reasoning, and the development of small open-source medical models. Our goal is to guide future research toward more trustworthy, transparent, and clinically viable LLMs.
@@ A Comprehensive Survey on the Trustworthiness of Large Language Models in Healthcare
Manar Aljohani $\spadesuit$, Jun Hou $\spadesuit$, Sindhura Kommu $\spadesuit$, Xuan Wang $\spadesuit$ $\spadesuit$ Department of Computer Science, Virginia Tech, Blacksburg, VA, USA {manara, junh, sindhura, xuanw}@vt.edu
(a) Temporal Trends
(b) Distribution of Datasets
(c) Distribution of Models
Figure 1: Overview of research trends, dataset usage, and model types across key trustworthiness dimensions in healthcare LLM studies: (a) Temporal Trends in Trustworthiness Dimensions Addressed in Medical LLM Studies (2022–2025); (b) Distribution of Dataset Types Across Trustworthiness Dimensions in Healthcare LLM Studies; (c) Distribution of Model Types Across Trustworthiness Dimensions in Healthcare LLM Studies.
@@ 1 Introduction
Truthfulness, defined as "the accurate representation of information, facts, and results by an AI system" Huang et al. (2024), is critical in healthcare, as inaccuracies can lead to misdiagnoses or inappropriate treatment recommendations. Ensuring that generated information is both accurate and aligned with verified medical knowledge is essential. Additionally, privacy concerns arise from the risk of exposing sensitive patient data during model training and usage, potentially leading to breaches or violations of regulations such as HIPAA (Health Insurance Portability and Accountability Act) and GDPR (General Data Protection Regulation). Ensuring patient confidentiality while leveraging LLMs for diagnostics and treatment recommendations is a critical challenge. Safety, defined as “ensuring that LLMs do not answer questions that can harm patients or healthcare providers in healthcare settings” Han et al. (2024b), further underscores the necessity of implementing stringent safeguards to mitigate harm. Robustness refers to an LLM’s ability to consistently generate accurate, reliable, and unbiased outputs across diverse clinical scenarios while minimizing errors, hallucinations, and biases. It also encompasses the model’s resilience against adversarial attacks, ensuring that external manipulations do not compromise its integrity. A truly robust LLM in healthcare must demonstrate stability, reliability, and fairness, even when faced with noisy, ambiguous, or adversarial inputs. Similarly, fairness and bias must be addressed to prevent discriminatory patterns in model outputs, which could lead to unequal treatment recommendations and exacerbate healthcare disparities. Furthermore, the explainability of LLMs, which ensures that model outputs are interpretable and transparent, plays a vital role in fostering trust and allowing informed decision-making by healthcare professionals. Lack of transparency in model reasoning complicates clinical adoption and raises accountability concerns.
Clinical deployments of LLMs expose trust gaps across dimensions. Med-PaLM and Med-PaLM 2 show truthfulness and safety issues, with hallucinated that could misguide care Singhal et al. (2023). Integrating LLMs with EHRs in cloud settings risks HIPAA/GDPR violations, prompting on-prem deployment and stronger de-identification Jonnagaddala and Wong (2025). Robustness remains problematic; frameworks like MEDIC and CREOLA assess hallucination severity and clinical safety Kanithi et al. (2024a); Asgari et al. (2025). Fairness issues persist, with studies showing that LLMs can perpetuate racial biases in medical recommendations Pfohl et al. (2024). Finally, explainability challenges were evident in AMIE—a conversational diagnostic agent evaluated in OSCE-style clinical exams—demonstrated strong diagnostic reasoning but lacked transparency compared to human doctors Tu et al. (2025).
@@
Tackling these challenges is essential for the trustworthy and ethical implementation of LLMs in healthcare. Recently, researchers have begun developing benchmarks and evaluation frameworks to systematically assess the trustworthiness of LLMs Huang et al. (2024). The trustworthiness of LLMs in healthcare is gaining increasing attention due to its significant social impact. However, there is currently no systematic review that provides a comprehensive understanding and future insights into this area. To bridge this gap, we present a comprehensive survey that explores these trust-related dimensions in detail, reviewing existing benchmarks and methodologies aimed at improving the trustworthiness of LLMs in healthcare.
@@ 2 Datasets, Models, and Tasks
2.1 Inclusion & Exclusion Criteria
We initiated our survey with a comprehensive literature search targeting studies on the trustworthiness of LLMs in healthcare. Our search strategy employed diverse keyword combinations and was directed toward top-tier conferences and journals, prioritizing publications from 2022 onward. Detailed inclusion and exclusion criteria are provided in Appendix A. Fig 1(a) illustrates how the number of papers addressing each key trustworthiness dimension in healthcare LLMs has changed over time from 2022 to 2025. From Figure 1(a), interest in trustworthiness dimensions peaked in 2024, particularly for Fairness and Bias (16 papers) and Privacy (11 papers), reflecting a strong recent push toward ethical and secure AI in healthcare. Truthfulness and Explainability maintained steady growth through 2023 and 2024. These trends suggest a rising concern with fairness and privacy in recent years, possibly driven by real-world deployment risks and regulatory pressure.
{forest}
forked edges, for tree= grow=east, reversed=true, anchor=base west, parent anchor=east, child anchor=west, base=center, font=, rectangle, draw=hidden-draw, rounded corners, align=left, text centered, minimum width=4em, edge+=darkgray, line width=1pt, s sep=3pt, inner xsep=2pt, inner ysep=3pt, line width=0.8pt, ver/.style=rotate=90, child anchor=north, parent anchor=south, anchor=center, , where level=1text width=14em,font=,, where level=2text width=18em,font=,, where level=3text width=18em,font=,, [ Trustworthiness of LLMs in Healthcare, ver [ Truthfulness, fill=blue!10 [ Benchmarks, fill=blue!10 [ Med-HALT Pal et al. (2023), PubHealthTab Akhtar et al. (2022), HEALTHVER Sarrouti et al. (2021), leaf, text width=66em ] ] [ Mitigation Methods, fill=blue!10 [ Self Reflection Ji et al. (2023), MEDAL Li et al. (2024), Faithful Reasoning Tan et al. (2024), HEALTHVER Sarrouti et al. (2021),
CRITIC Gou et al. (2024), SEND Mohammadzadeh et al. (2024), leaf, text width=66em ] ] [ Evaluation and Detection Methods, fill=blue!10 [ Med-HALT Pal et al. (2023), Med-HVL Yan et al. (2024), Semantic Entropy Farquhar et al. (2024),
SEPs Han et al. (2024a), Faithful Reasoning Tan et al. (2024), PubHealthTab Akhtar et al. (2022), HEALTHVER Sarrouti et al. (2021),
CRITIC Gou et al. (2024), Cross-Examination Cohen et al. (2023), MAD Smit et al. (2023), leaf, text width=66em ] ] ] [ Privacy, fill=blue!10 [ Benchmarks, fill=blue!10 [ SecureSQL Song et al. (2024), leaf, text width=66em ] ] [ Enhance Methods, fill=blue!10 [ Federated Learning Zhao et al. (2024a),
Differential Privacy Singh et al. (2024), De-identification Liu et al. (2023b), Mitigating Memorization Yang et al. (2024a), APNEAP Wu et al. (2024), leaf, text width=66em ] ] [ Evaluation Methods, fill=blue!10 [ SecureSQL Song et al. (2024), Memorize Fine-tuning Data Yang et al. (2024a), clinical Note De-identification Altalla’ et al. (2025),
Memorization Yang et al. (2024a), Textual Data Sanitization Xin et al. (2024), leaf, text width=66em ] ] ] [ Safety, fill=blue!10 [ Benchmarks, fill=blue!10 [ Med-harm Han et al. (2024c), Medsafetybench Han et al. (2024b), leaf, text width=66em ] ] [ Enhance Methods, fill=blue!10 [ UNIWIZ Das and Srihari (2024), Data-Poisoning Attack Han et al. (2024e), leaf, text width=66em ] ] [ Evaluation Methods, fill=blue!10 [ Med-harm Han et al. (2024c), Medsafetybench Han et al. (2024b), Misinformation Attacks Han et al. (2024d), MEDIC Kanithi et al. (2024b),
GLiR Attack Leemann et al. (2024), Data-Poisoning Attack Han et al. (2024e), leaf, text width=66em ] ] ] [ Robustness, fill=blue!10 [ Benchmarks, fill=blue!10 [ Detecting Anomalies Rahman et al. (2024), RobustQA Han et al. (2023), RABBITS Gallifant et al. (2024), leaf, text width=66em ] ] [ Enhance Methods, fill=blue!10 [ LLM-TTA O’Brien et al. (2024), Detecting Anomalies Rahman et al. (2024), Secure Your Model Tang et al. (2024), MEDSAGE Binici et al. (2025),
Out-of-Context Prompting Cotta and Maddison (2024), leaf, text width=66em ] ] [ Evaluation Methods, fill=blue!10 [ Stumbling Blocks Wang et al. (2024), Detecting Anomalies Rahman et al. (2024),
Instruction Phrasings Ceballos-Arroyo et al. (2024), RobustQA Han et al. (2023), RABBITS Gallifant et al. (2024), leaf, text width=66em ] ] ] [ Fairness and Bias, fill=blue!10 [ Benchmarks, fill=blue!10 [ BiasMedQA Schmidgall et al. (2024), EquityMedQA Pfohl et al. (2024), Superficial Fairness Alignment Wei et al. (2024), FairMedFM Jin et al. (2024), leaf, text width=66em ] ] [ Mitigation Methods, fill=blue!10 [ BiasMedQA Schmidgall et al. (2024), Reinforcement Learning with Clinician Feedback Zack et al. (2024), Instruction Fine-tuning Singhal et al. (2023),
Hurtful Words Zhang et al. (2020), Mitigate Cognitive Biases Ke et al. (2024), CI4MRC Zhu et al. (2023), Bias of Disease Prediction Zhao et al. (2024b),
Racial and LGBTQ+ Biases Xie et al. (2024), Out-of-Context Prompting Cotta and Maddison (2024), Attribute Neutral Modeling Hu et al. (2024),
Personalized Alignment Techniques Kirk et al. (2024), Evaluating Biases in Context-Dependent Patel et al. (2024), leaf, text width=66em ] ] [ Evaluation and Detection Methods, fill=blue!10 [ Evaluation Study Zack et al. (2024), BiasMedQA Schmidgall et al. (2024), Hurtful Words Zhang et al. (2020), Race-based Medicine Omiye et al. (2023),
Detect Debunked Stereotypes Swaminathan et al. (2024), EquityMedQA Pfohl et al. (2024), Superficial Fairness Alignment Wei et al. (2024),
Examines Biased AI Adam et al. (2022), Identify Biases Yang et al. (2024b), Quantifying Cognitive Biases Lin and Ng (2023),
Biases in Biomedical MLM Kim et al. (2023), Bias of Disease Prediction Zhao et al. (2024b), Racial and LGBTQ+ Biases Xie et al. (2024),
FairMedFM Jin et al. (2024), leaf, text width=66em ] ] ] [ Explanability, fill=blue!10 [ Benchmarks, fill=blue!10 [ FaReBio Fang et al. (2024), Pathway2Text Yang et al. (2022), leaf, text width=66em ] ] [ Enhance Methods, fill=blue!10 [ Knowledge Graphs Shariatmadari et al. (2024), Medical Imaging Explainability Ghosh et al. (2023), MedExQA Kim et al. (2024),
Retrieval and Reasoning on KGs Ji et al. (2024), DDCoT Zheng et al. (2023), A ChatGPT Aided Explainable Framework Liu et al. (2023a),
Medical Concept-Driven Attention Wang et al. (2022), FaReBio Fang et al. (2024),
LLM-GCE He et al. (2024), kNN-Graph2Text Yang et al. (2022), RAG-IM Mahbub et al. (2024), MedThink Gai et al. (2025), leaf, text width=66em ] ] ] ]
2.2 Datasets
@@
The datasets used in studies of trust in LLMs for healthcare are categorized by the dimensions of trustworthiness they address in Appendix B, where we highlight key details such as data type, content, task, and dimensions of trustworthiness. The content of each dataset specifies its composition, while the task refers to the main problem to be solved for which the dataset is utilized. The data type varies across studies and includes web-scraped data, curated domain-specific datasets, public text corpora, synthetic data, real-world data, and private datasets, providing a comprehensive overview of their relevance to healthcare applications.
Figure 1(b) shows the number of studies using three major dataset types—Med-QA (blue), Med-Gen (orange), and Med-IE (green)—in relation to six trustworthiness dimensions: Truthfulness, Privacy, Safety, Robustness, Fairness and Bias, and Explainability. Figure 1(b) shows how three major dataset types—Med-QA, Med-Gen, and Med-IE—are used across six trust dimensions. Truthfulness is most studied with both Med-QA and Med-Gen. Med-QA is also common in fairness and explainability, while Med-Gen contributes to safety and privacy. Med-IE, though less used overall, is more prominent in robustness and explainability. This highlights the dominance of Med-QA and Med-Gen, with Med-IE offering value in specific areas of trustworthiness.
2.3 Models
The models assessed in studies on trust in LLMs for the healthcare domain are outlined, along with their trustworthiness dimensions, in Appendix C, where we summarized key details such as the model name, release year, openness, architecture, task, and the institution responsible for its development. Figure 1(c) illustrates the proportions of different model types—open-source, closed-source, and architectures including encoder-only, decoder-only, and encoder-decoder—used in research addressing various trustworthiness aspects of LLMs in healthcare: Explainability, Fairness and Bias, Robustness, Safety, Privacy, and Truthfulness. From Figure 1(c), it is clear that Decoder-only and Open-source models are the most commonly used across all trustworthiness dimensions—especially in robustness, explainability, and truthfulness—highlighting their accessibility and alignment with generative tasks. Closed-source models appear more in fairness and privacy studies, while Encoder-only and Encoder-decoder models are used less frequently, mostly in fairness and truthfulness evaluations.
2.4 Tasks
The tasks covered various primary focuses of LLMs in healthcare. Inspired from the survey by Liu et al. (2024a), these tasks include:
@@ 3 Trustworthiness of LLMs in Healthcare
We examine the challenges related to the trustworthiness of LLMs in healthcare, outlining key strategies for identifying and mitigating these concerns. From our literature review screening, we identified truthfulness, privacy, safety, robustness, fairness and bias, and explainability as key trustworthiness dimensions of LLMs as highlighted in TrustLLM Huang et al. (2024), particularly in healthcare. Figure 2 provides a summary of the recent research on trust in LLMs for healthcare across key dimensions of trustworthiness.